During a non-flow process 120 Btu are removed as heat from each pound of working substance while the internal energy decreases 85.5 Btu /lb. Calculate the work involved in the process in ft·lbf/lb
1
Expert's answer
2020-03-16T13:07:10-0400

By the law of thermodynamics

"\Delta U=Q-W"

"W=Q-\Delta U"

As 120 Btu of heat is removed from the system which means "Q=-120\ \\"

and "\Delta U=-85.5 Btu"

"W=-120-(-85.5)=-120+85.5=-34.5Btu"